OpenStudy (fools101):

I Don't know How to do this PLEASE HELP!!! Given f(x) = (x - 3)2 + 2 and g(x) = mathml equation Show they are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (fools101):



OpenStudy (anonymous):

same here i neep help on it too we have 30 questions then 49 more omg so much math

OpenStudy (fools101):

i know it Driving me Crazy @.@

OpenStudy (fools101):

Can u help me it Driving me Crazy i'm sooo Confuse!! please!

OpenStudy (anonymous):

plug em in

OpenStudy (anonymous):

it so got damn hard

OpenStudy (anonymous):

im srry i passed math but i dont get this

OpenStudy (anonymous):

f(g(x)) = f ( sqrt (x-2) + 3 ) = ( [sqrt (x-2) + 3 ] - 3 ) +2

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

f(g(x)) = f ( sqrt (x-2) + 3 ) = ( [sqrt (x-2) + 3 ] - 3 )^2 +2 = x-2 +2 = x

OpenStudy (anonymous):

ok that works,

OpenStudy (fools101):

ohh thank u so much i have Struggle with it thank u sooo much Barry2 thanks for trying Crank dat History

OpenStudy (anonymous):

i got the answer

OpenStudy (anonymous):

when we square a number we use ^, like x^2